use two unit multipliers to convert 36 feet to centimeters. (go from feet to inches to centimeters.
Answer by ikleyn(52785)

.

36 feet = 36*12 inches = 36*12*2.54 centimeters = 1097.28 centimeters.    ANSWER

Solved.